OpenStudy (anonymous):

Derivative of cos x^2 by abinitio method.......?

OpenStudy (anonymous):

@sakh @ParthKohli @ParthKohli

OpenStudy (rogue):

I don't know what the abinitio method is, but you can differentiate this by using the chain rule.\[\frac {d}{dx} \left[ \cos^2 x \right] = 2 \cos x * \frac {d}{dx} \left[ \cos x \right] = -2 \cos x \sin x = -\sin 2x\]